Grade 12th passMechanics

a 500 kg boat of length 9 m and is floating without motion in still water. a man of 100 kg is at one end and if he runs to other end of boat and stops, the displacement of boat is

Vidit goel
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Let the boat be displaced by x istance when the man runs from one end to other end.
Therefore from momenta conservation principle in this disaplcemnet form we get,
M*x = m(l-x)
on solving we get,
x = ml/(M+m) = > 100*9/(100+500) = > 900/600 = > 9/6 => 3/2 = 1.5 meter.